The P/E ratio of the S&P 500 should be 18 based on historical analysis and your projection of the future. If the earnings of the S&P 500 are currently $40 then the S&P 500 should be at:
Answer #1

If projection of P/E is 18 then

P/E = 18

For earning = $40

P/40 = 18 or P = 40*18

P = $720

So S&P 500 should be at 720